### v3

# Optimizing Sparse Convolution on GPUs with CUDA for 3D Point Cloud Processing in Embedded Systems

Abstract
The increasing utilisation of LiDAR and 3D sensors has made the analysis of 3D point clouds indispensable in various applications, including object detection and segmentation. In contrast to photos, point clouds possess distinctive computing issues due to their sparse nature and absence of a regular grid. Traditional Convolutional Neural Networks (CNNs) that are designed to perform well on dense data are not suitable for this particular task. The utilisation of sparse neural networks specifically engineered for the efficient handling of sparse data is imperative in this context. Nevertheless, the utilisation of neural networks on Graphics Processing Units (GPUs), particularly with the CUDA framework, presents challenges due to the irregular nature of point cloud data and the requirement for optimised memory access patterns. The present study focuses on enhancing the efficiency of sparse convolution operators for 3D point clouds on GPUs through the utilisation of CUDA technology. This paper presents a novel approach that combines the theoretical benefits of sparse neural networks with efficient GPU-based implementations. By doing so, it provides valuable insights and techniques for effectively using 3D point cloud analysis, hence enhancing the capabilities of object detection and segmentation in various fields.
Index Terms:
Sparse Convolution, CUDA, Point CloudI Introduction
In recent years, there has been a significant increase in the utilization of deep learning methods, particularly convolutional neural networks (CNNs) [articlecnn], which have emerged as the dominant approach in various domains that involve structured grid data, such as picture analysis and processing. Nevertheless, the exponential growth in the utilization of LiDAR and 3D sensors across many domains has resulted in an increased need for the analysis of 3D point clouds. The utilization of 3D point clouds is crucial in various applications, including object recognition and segmentation, as they offer a spatial depiction of things within a three-dimensional environment. In contrast to photos, point clouds exhibit sparsity and lack a regular grid, hence posing distinct processing and computational issues.
The examination of 3D point cloud data has initiated a novel period of progress in the field of computer vision, enabling a wide range of applications that use its abundant spatial information. One of the most popular application of point cloud data is automated driving [fernandes2021point] and deep learning can play role in various task on point cloud such as 3D shape classification, 3D object detection and tracking, 3D point cloud segmentation and 3D point cloud registration [9127813]. Qi et al. introduced a significant contribution in the form of ”PointNet” [qi2017pointnet], which marked a notable departure from existing approaches. This study presented a deep learning framework that was tailored to handle point sets, with a particular emphasis on addressing challenges related to 3D classification and segmentation tasks. In addition, Wang et al. introduced a real-time semantic segmentation method named ”PointSeg” [wang2018pointseg] that focuses on utilizing 3D LiDAR point clouds for this purpose. The aforementioned advancements highlight the extensive capabilities of 3D point cloud analysis in diverse practical situations, ranging from autonomous navigation to virtual reality.
Point clouds, distinct from traditional image data due to their sparsity, present challenges in applying standard CNN architectures, which are better suited for dense data. This sparsity, a result of the data collection methodology and inherent information gaps in 3D space, necessitates the adoption of sparse neural networks. These networks are tailored to efficiently process sparse data while reducing computational demands. Sparse Convolution Neural Network (SCNN) is regarded as solution to applied deep learning on sparse data like point cloud. However, implementing them, particularly on GPUs, is complex due to the irregular data structures and the need for optimized memory access patterns.
To bridge the gap between the computational intensity of SCNNs and the limited processing capabilities of edge computing devices, current developments in autonomous driving and 3D hand scanning technologies underscore the necessity for real-time processing of 3D point cloud data. These applications often operate in scenarios where rapid data processing is critical for decision-making and user interaction, yet they are constrained by the available computational resources. This dichotomy highlights the importance of efficient, real-time 3D data processing solutions that can operate within the confines of edge computing devices. NVIDIA’s Jetson platform, equipped with CUDA technology, emerges as a suitable candidate for such tasks, offering a balance between computational power and energy efficiency, making it ideal for handling the demands of real-time, sparse data processing in applications like autonomous vehicles and advanced scanning devices. Recently, NVIDIA’s introduction of the TensorRT [vanholder2016efficient] inference framework has shown promise in addressing these issues. TensorRT, which allows custom plugin development, enhances the feasibility of deploying neural networks, specifically those utilizing sparse convolution, on embedded systems.
This research investigates the challenges of implement sparse convolution efficiently utilising GPUs on Jetson Platform with CUDA, to improve the speed of performing inference on sparse convolution operators for 3D point clouds. The main goal is to establish a correlation between the potential advantages of sparse neural networks and their practical application in real-life scenarios, and we will provide CUDA solutions for three categories of sparse convolution operators. This paper is structured as follows: Section II will provide a comparison between sparse convolution and traditional convolution to highlight their distinct characteristics. Additionally, it will present the current research on this topic. Section III presents a highly efficient method for implementing sparse convolution on CUDA and section IV outlines our discussions and conclusions.
II Background and Motivation
II-A Structure and Represention for Point Cloud Data
Point cloud data, a fundamental representation in 3D data analysis, consists of data points within a three-dimensional coordinate system, typically generated by 3D scanning technologies. These points, representing the external surfaces of objects, are characterized by their density, distribution, and precision. Traditional point cloud representation primarily involves raw coordinates, sometimes enriched with attributes like color, normal, and reflectance. However, the unstructured nature of raw point clouds poses challenges in computer vision and graphic because they are inherently complex and irregular, consist of a massive number of points and contain noise and outliers. Voxelization, the conversion of continuous point cloud data into a discrete 3D grid, addresses these challenges. Each voxel in the grid represents the presence or characteristics of points within its boundaries, thereby regularizing the data for computational efficiency in storage, compression, and processing. [XU2021103675]
When they need to be processed by neural network, for a point cloud and its voxelized data, the sparse tensor might encapsulate the actual 3D coordinates of points, alongside any associated attributes such as color, intensity, or normal. In typical sparse convolution implementations, data are stored in memory using specialized structures to efficiently represent and process the sparsity. The indices structure holds the coordinates of non-zero points in a sparse tensor, which corresponds to voxels or points in a point cloud. This structure allows the algorithm to iterate only over the meaningful data points, significantly reducing the computational load. For each batch in a dataset, the indices are stored with their batch ID, allowing for parallel processing of multiple data instances. The data structure for this work [NEURIPS2019_57370345] [article3dvoxel]